OpenStudy (anonymous):

Use Cramer's rule to find the solution of the system of equations. Write your answer as an ordered pair. 3x+y=5 2x-3y=18

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

get the determinant first 3 1 2 -3 the determinant is -9-2=-11. then just replace the variable you want to solve for with the answer like in the other problem. so it would be 5 1 18 -3 to find x which wold be -33. and then you divide that by the det, which is -11 so x=-33/-11=3 x=3 3 5 2 18 which =44 44/-11=4 y=-4

OpenStudy (anonymous):

Thank you so much! I really appreciate it :)

OpenStudy (anonymous):

Welcome:] Can i have medal

OpenStudy (wolf1728):

I think I'll add a little more to that so it is easier to see Cramer's rule at work. 3x+y=5 2x-3y=18 a = 3 b =1 c = 2 d = -3 e = 5 f = 18 We first have to find a denominator determinant dn = (a • d) - (c • b) dn = (3•-3) -(2•1) dn = -11 Next we have to figure the x determinant and divide it by dn x = [(e • d) - (f • b)] ÷ dn x = (5*-3) -(18 •1) / -11 x = -15 -18 / -11 x = -33 / -11 x = 3 Now for the y determinant y = [(a • f) -(c • e)] ÷ dn y = (3 • 18) -(2 • 5) / -11 y = 54 -10 / -11 y = 44 / -11 y = -4 Source: http://www.1728.org/cramer.htm